OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin^2x = sin x I watched a live lesson, but I'm still completely stuck... lol

OpenStudy (mayankdevnani):

\[\large \bf 2\sin^2x=sinx\] \[\large \bf 2\sin^2x-sinx=0\] Take `sinx ` common, \[\large \bf sinx(2sinx-1)=0\] so either sinx=0 OR (2sinx-1)=0

OpenStudy (mayankdevnani):

solve these 2 equations and find solutions in the interval [0,2pi)

OpenStudy (mayankdevnani):

hope you understand @zeeke

OpenStudy (anonymous):

@mayankdevnani I understand up until the very last part.... How would I solve those? Substitute in a trigonometric identity...? I'm slightly confused. These are the answer choices btw, I mistakenly left them out. x = π/3, 2π/3 x = π/2, 3π/2, π/3, 2 π / 3. x = 0, π, π/six, 5π/6 x = π/6, 5π/6

OpenStudy (mayankdevnani):

what you think ?

OpenStudy (mayankdevnani):

i think the most appropriate answer would be option C but to correct option C we have to change the interval \[\large \bf OLD~INTERVAL \rightarrow [0,2 \pi)\] \[\large \bf NEW~INTERVAL \rightarrow [0,2 \pi]\] then answer would be option C otherwise option D is answer

OpenStudy (mayankdevnani):

understood ? @zeeke

OpenStudy (anonymous):

@mayankdevnani I'm so sorry, but I'm honestly still kinda lost. Where exactly do these intervals come into play? The lesson I took seemed to teach everything BESIDES these interval sort of problems, so I'm completely taken aback by these questions...

OpenStudy (anonymous):

I'm just not quite sure of what to do after I know sinx=0 and (2sinx-1)=0

OpenStudy (anonymous):

I'm just not quite sure of what to do after I know sinx=0 and (2sinx-1)=0